OpenStudy (anonymous):

Compute the following (i)^(-n) please show work so I can understand how to get the answer. Thank you

OpenStudy (anonymous):

\[b^{-n}=\frac{1}{b^n}\]

OpenStudy (anonymous):

is \(i=\sqrt{-1}\) in this case?

OpenStudy (anonymous):

no, I dont think so

OpenStudy (anonymous):

then by definition \[i^{-n}=\frac{1}{i^n}\]there is no other work to be done

OpenStudy (anonymous):

We have to be able to express it in a+ib form